OpenStudy (anonymous):

Write two fractions that are equivalent. 1.) 6/12 2.) 14/16 3.) 16/20 4.) 7/10

OpenStudy (anonymous):

hey

hero (hero):

multiply each fraction by 2/2, then multiply each fraction by 3/3

OpenStudy (anonymous):

ty hero.. lol stole the words outta my mouth.

hero (hero):

Sorry. I don't mean to steal anything :P

OpenStudy (anonymous):

im not mad im used to u always beating me to math. lol

OpenStudy (anonymous):

Thank you very much Hero. Medal given for your intelligence in Mathematics.

hero (hero):

:D

OpenStudy (anonymous):

Hey hero... ugh.. help me.. please?

hero (hero):

(6x2)/(12x2) = 12/24

OpenStudy (anonymous):

Would that be one way?

OpenStudy (anonymous):

And then to get the second would I take (6x3)/12x3) = 18/36?

hero (hero):

Yes, you are doing it correctly

OpenStudy (pokemon23):

Hero can you help the poor fellow? I understand his pain

OpenStudy (pokemon23):

thanks

OpenStudy (anonymous):

So I would do the same for 14/16 I could multiply it like (14x2)/16x2) = 28/32?

OpenStudy (pokemon23):

you multiply right ?

hero (hero):

Yes, you have it miraj. It's not as much of a mystery as you think. You might want to use a calculator to check your answers.

OpenStudy (anonymous):

Thanks Hero just want to double check. I think too much and worry a lot. Pokemon please read my biography on my profile page. I am a Sophomore struggling in Mathematics only.. Please have their weaknesses and Math is my weakness. All my friends are in Algebra 2 or Pre Calculus but I'm the one left behind. I do find this easy but also difficult to understand as well. I am absolutely the worst when it comes to doing fractions.

OpenStudy (pokemon23):

I understand

OpenStudy (pokemon23):

you should see my question I put up your the not only person that struggles this year I'm doing better than math than expected

OpenStudy (pokemon23):

usually in 8th grade I did poorly on math

OpenStudy (pokemon23):

and right now I have to take a two year course in Algebra

hero (hero):

Sounds like you just had a bad teacher @ Miraj

OpenStudy (pokemon23):

i feel his pain :'( my teacher was so lazy all he did is say go look it up on the internet

OpenStudy (anonymous):

Indeed Hero I did. I had to tell the principle that he swore non stop to the good students and treated the bad students like his friends. He slammed the door shut so loud that you can hear it across the other side of the school whenever he got mad and he would get really mad when you came up to him to ask for more help. I'm gonna let this out there but I cried in class one day as a man because I felt such pain that he didn't help me and I wanted to know how to do the problems. I needed to get an A in the class and he ended up failing me on my final exam.

OpenStudy (pokemon23):

I know your pain my old math teacher screamed at face and made fun of my classmates and made one kid cry so badly everyone was shock

hero (hero):

I've never been in your situation so I don't know what I would do. Math was always easy for me and I had mostly women teachers.

OpenStudy (pokemon23):

lucky women teachers are nicer

OpenStudy (pokemon23):

right now I have a women teacher and I have bee doing so well in her class she doesn't pick me do questions on the board because she knows I'm smart